A hydraulic pump supplies 30 litres/second. How long will it take to fill a 10 x 10 m tank that is 1 m deep?
Answer by xinxin(76)

The question indicates that the tank is a rectangle prism (length 10m, width 10m, height 1 m). The volume = 10*10*1 = 100 m^3. Next we need to covert from one unit to the other, either from liter to m^3 or from m^3 to liter (the second way would be easier because it has only integers). Since 1 m^3 = 1000 liters, the volume of the tank = 100*1000 = 10^5 liters, so the time it takes to fill the tank = 10^5/30 = 3333 seconds.